Hi were would I insert the Parrentheses to get an answer of 8
16 divided by 2 X 3 + 2 + 6

Answer by stanbon(75887)

Hi were would I insert the Parrentheses to get an answer of 8
16 divided by 2 X 3 + 2 + 6
-------
16 /[2*3+2]+6 = 16/8 + 6 = 2+6 = 8
-----------------------------------
Cheers,
Stan H.